# Asymptotics for conformal inference

### v1

@@ 5.2.2 Joint convergence and application to the $\mathrm{FDP}$ and $\mathrm{TDP}$
Following Neuvial, (2008), $G\mapsto\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)$ and $G\mapsto\mathrm{FDP}_{m}^{(n)}(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G))$ are Hadamard differentiable at $G_{{\tiny\mbox{mixt}}}$ provided that $G_{{\tiny\mbox{mixt}}}$ is concave and differentiable and that $\alpha>[G_{{\tiny\mbox{mixt}}}^{\prime}(0^{+})]^{-1}$. The following result complete the picture by studying the convergence of ${\widehat{G}}_{m}^{(n)}$.
@@
$\sqrt{\tau_{n,m}}\begin{pmatrix}{\widehat{{G}}}_{m,{0}}^{(n)}-I\\ {\widehat{{G}}}_{m,{1}}^{(n)}-G\end{pmatrix}\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\begin{pmatrix}\frac{\sigma}{\sqrt{\pi_{0}}}\mathbb{U}+\sqrt{1-\sigma^{2}}\mathbb{W}\\ \frac{\sigma}{\sqrt{1-\pi_{0}}}\mathbb{V}_{G}+G^{\prime}\sqrt{1-\sigma^{2}}\mathbb{W}\end{pmatrix}\eqqcolon\begin{pmatrix}\mathbb{Z}_{0}\\ \mathbb{Z}_{1}\end{pmatrix}\text{ on $\left[D(0,1)\right]^{2}$},$
@@
$\sqrt{\tau_{n,m}}\left({\widehat{G}}_{m}^{(n)}-G_{{\tiny\mbox{mixt}}}\right)\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}$ $\sqrt{\pi_{0}\sigma^{2}}\mathbb{U}+\sqrt{\left(1-\pi_{0}\right)\sigma^{2}}\mathbb{V}_{G}+G_{{\tiny\mbox{mixt}}}^{\prime}\sqrt{1-\sigma^{2}}\mathbb{W}$
$=\pi_{0}\mathbb{Z}_{0}+\left(1-\pi_{0}\right)\mathbb{Z}_{1}\eqqcolon\mathbb{Z}\text{ on $D(0,1)$}.$
@@
By using the functional delta method theorem with the BH functionals described above (see Neuvial, (2008), supplementary files from Delattre and Roquain, (2016) and Lemma S.2.2 from Kluger and Owen, (2024)), we obtain the following result.
@@
$\sqrt{\tau_{n,m}}\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}({\widehat{G}}_{m}^{(n)})-\mathcal{T}_{\alpha}\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\frac{1}{\frac{1}{\alpha}-G_{{\tiny\mbox{mixt}}}^{\prime}(\mathcal{T}_{\alpha})}\mathbb{Z}({\mathcal{T}_{\alpha}});$
$\sqrt{\tau_{n,m}}\left(\mathrm{FDP}_{m}^{(n)}\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}({\widehat{G}}_{m}^{(n)})\right)-\pi_{0}\alpha\frac{}{}\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\frac{\pi_{0}}{G_{{\tiny\mbox{mixt}}}(\mathcal{T}_{\alpha})}\mathbb{Z}_{0}({\mathcal{T}_{\alpha}});$
$\sqrt{\tau_{n,m}}\left(\mathrm{TDP}_{m}^{(n)}\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}({\widehat{G}}_{m}^{(n)})\right)-G(\mathcal{T}_{\alpha})\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\frac{G^{\prime}(\mathcal{T}_{\alpha})}{\frac{1}{\alpha}-G_{{\tiny\mbox{mixt}}}^{\prime}(\mathcal{T}_{\alpha})}\mathbb{Z}({\mathcal{T}_{\alpha}})+\mathbb{Z}_{1}(\mathcal{T}_{\alpha}),$
@@
Proposition 5.3 is proved in Section B.3. As a sanity check, we see that Proposition 5.3 reduces to the result of Neuvial, (2008) when $n/m$ tends to infinity ($\sigma^{2}=1$).
@@
$\begin{pmatrix}\mathbb{Z}_{0}\\ \mathbb{Z}_{1}\\ \mathbb{Z}\end{pmatrix}=\begin{pmatrix}\frac{\sigma}{\sqrt{\pi_{0}}}\mathbb{U}+\sqrt{1-\sigma^{2}}\mathbb{W}\\ \frac{\sigma}{\sqrt{1-\pi_{0}}}\mathbb{V}_{G}+G^{\prime}\sqrt{1-\sigma^{2}}\mathbb{W}\\ \sqrt{\pi_{0}\sigma^{2}}\mathbb{U}+\sqrt{\left(1-\pi_{0}\right)\sigma^{2}}\mathbb{V}_{G}+G_{{\tiny\mbox{mixt}}}^{\prime}\sqrt{1-\sigma^{2}}\mathbb{W}\end{pmatrix},$
@@
Proving Theorem 4.2 is analogue to the proof of Theorem 4.1 above, but starting from weighted processes. For short, the full description of the proof is postponed to Section C.3. Therein, Propositions C.3 and C.4 are the analogues of Propositions 5.2 and 5.3, respectively.
@@ 6 Conclusion
In this paper we obtained the exact asymptotic distribution of $\mathrm{FCP}$ and $\mathrm{FDP}$ for conformal inference methods when both the sizes of the calibration sample and test sample grow simultaneously. Our theory covered both the prediction and novelty detection settings, including a potential distribution shift. Our results quantified exactly how the covariance process is affected by the dependence inherent to the conformal settings, that use the same calibration sample for all the test examples. First, we proved that the convergence rate $\tau_{n,m}$ can be largely deteriorated when $n/m$ vanishes to zero (that is, $\sigma=0$). Otherwise, when $n$ and $m$ are of the same order, the convergence rate is the usual one ($m^{1/2}$), but the asymptotic covariance is affected. Nevertheless, when $n/m$ tends to infinity, the convergence is strictly the same as in the usual independent case. Interestingly, our results can be used to calibrate easily and accurately quantiles for controlling an error amount when performing conformal inference with large $n$ and $m$. We also quantified the effects of doing conformal inference while there is a distribution shift between the calibration sample and the test sample. We exhibits how this distribution shift acts on the asymptotic behaviour of the $\mathrm{FCP}$, by changing the mean and the variance, and how the correction with weighted conformal $p$-values impacts the asymptotic variance.
While our work paves the way for studying asymptotic convergences in conformal inferences, it left some open directions for future research. For instance, in case of a distribution shift, the oracle weight function is mostly unknown, and is often estimated (Jin and Candès,, 2023). Finding the exact asymptotic distribution for the processes using estimated weights is a very interesting and challenging problem for future investigations.
Acknowledgements
I am grateful to Étienne Roquain for his precious advices and reviews throughout this work. I wish to thank Gilles Blanchard and Sylvain Delattre for their comments, and Magalie Fromont for fruitful discussions. I acknowledge the Emergence project MARS of Sorbonne Université.
@@ Appendix A Standardisation lemma
with $S^{\prime}_{k}=F_{\tiny\mbox{cal}}(S_{k})$ which are iid uniform and $T^{\prime}_{i}=F_{\tiny\mbox{cal}}(T_{i})$ which are iid $\sim{F_{\tiny\mbox{test,cal}}}$ (or either $\sim{F_{\tiny\mbox{$0$,cal}}}$ under the null or $\sim{F_{\tiny\mbox{test,cal}}}$ under the alternative, in the novelty detection setting). ∎
@@ Appendix B Proofs of auxiliary results
In this section, we prove Proposition 5.1, Propositions 5.2 and 5.3. Proofs for the weighted case are given in Section C.
@@
$\sqrt{m}\left({\widehat{F}}_{m,\tiny\mbox{test}}-F_{\tiny\mbox{test}}\right)\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\mathbb{U}\circ F_{\tiny\mbox{test}}\text{ on $D(\mathbb{R})$,}$
@@
$\sqrt{n}\left({\widehat{F}}_{\tiny\mbox{cal}}^{(n)}-F_{\tiny\mbox{cal}}\right)\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\mathbb{V}\circ F_{\tiny\mbox{cal}}\text{ on $D(\mathbb{R})$,}$
@@
$\sqrt{n}\left({({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}-F_{\tiny\mbox{cal}}^{-1}\right)\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}({F_{\tiny\mbox{cal}}^{-1}})^{\prime}\mathbb{V}\text{ on $D(0,1)$,}$ (43)
@@
$\sqrt{n}\left(F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}-F_{\tiny\mbox{test}}\circ F_{\tiny\mbox{cal}}^{-1}\right)\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}F_{\tiny\mbox{test}}^{\prime}\circ F_{\tiny\mbox{cal}}^{-1}({F_{\tiny\mbox{cal}}^{-1}})^{\prime}\mathbb{V}=\left(F_{\tiny\mbox{test}}\circ F_{\tiny\mbox{cal}}^{-1}\right)^{\prime}\mathbb{V}\text{ on $D(0,1)$.}$
Since for all $t\in(0,1)$, $G^{\prime}(t)=\left(F_{\tiny\mbox{test}}\circ F_{\tiny\mbox{cal}}^{-1}\right)^{\prime}(1-t)$ and since $\mathbb{V}$ is a standard Brownian bridge if and only if $(\mathbb{V}_{t})_{t}=(\mathbb{V}_{1-t})_{t}$ is a standard Brownian bridge we obtain the second term in (33) .
@@
$\begin{pmatrix}\sqrt{m}\left[{\widehat{F}}_{m,\tiny\mbox{test}}-F_{\tiny\mbox{test}}\right]\\ \sqrt{n}\left[F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ F_{\tiny\mbox{cal}}^{-1}(1-I)\right]\end{pmatrix}\overset{\mathcal{L}}{\rightarrow}\begin{pmatrix}\mathbb{U}\circ F_{\tiny\mbox{test}}\\ G^{\prime}\mathbb{V}\end{pmatrix}\text{ on $D(\mathbb{R})\times D(0,1)$},$
@@
$\begin{pmatrix}\sqrt{m}\left[{\widehat{F}}_{m,\tiny\mbox{test}}-F_{\tiny\mbox{test}}\right]\\ \sqrt{n}\left[F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ F_{\tiny\mbox{cal}}^{-1}(1-I)\right]\\ {({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)\end{pmatrix}\overset{\mathcal{L}}{\rightarrow}\begin{pmatrix}\mathbb{U}\circ F_{\tiny\mbox{test}}\\ G^{\prime}\mathbb{V}\\ F_{\tiny\mbox{cal}}^{-1}(1-I)\end{pmatrix}$
[…]
@@
$\begin{pmatrix}\sqrt{m}\left[{\widehat{F}}_{m,\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)\right]\\ \sqrt{n}\left[F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ F_{\tiny\mbox{cal}}^{-1}(1-I)\right]\end{pmatrix}\overset{\mathcal{L}}{\rightarrow}\begin{pmatrix}\mathbb{U}\left(F_{\tiny\mbox{test}}\circ F_{\tiny\mbox{cal}}^{-1}(1-I)\right)\\ G^{\prime}\mathbb{V}\end{pmatrix}$
[…]
Since $\mathbb{U}\left(F_{\tiny\mbox{test}}\circ F_{\tiny\mbox{cal}}^{-1}(1-I)\right)$ has the same distribution as $\mathbb{U}(G)$, we obtain (33).

### v3

@@ 5.2.2 Joint convergence and application to the $\mathrm{FDP}$ and $\mathrm{TDP}$
Following Neuvial (2008), $G\mapsto\mathcal{T}^{\mathrm{BH}_{\alpha}}(G)$ and $G\mapsto\mathrm{FDP}_{m}^{(n)}(\mathcal{T}^{\mathrm{BH}_{\alpha}}(G))$ are Hadamard differentiable at $G_{{\tiny\mbox{mixt}}}$ provided that $G_{{\tiny\mbox{mixt}}}$ is concave and differentiable and that $\alpha>[G_{{\tiny\mbox{mixt}}}^{\prime}(0^{+})]^{-1}$. The following result complete the picture by studying the convergence of ${\widehat{G}}_{m}^{(n)}$.
@@
$\sqrt{\tau_{n,m}}\begin{pmatrix}{\widehat{{G}}}_{m,{0}}^{(n)}-I\\ {\widehat{{G}}}_{m,{1}}^{(n)}-G\end{pmatrix}\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\begin{pmatrix}\frac{\sigma}{\sqrt{\pi_{0}}}\mathbb{U}+\sqrt{1-\sigma^{2}}\mathbb{W}\\ \frac{\sigma}{\sqrt{1-\pi_{0}}}\mathbb{V}_{G}+G^{\prime}\sqrt{1-\sigma^{2}}\mathbb{W}\end{pmatrix}\eqqcolon\begin{pmatrix}\mathbb{Z}_{0}\\ \mathbb{Z}_{1}\end{pmatrix}\text{ on $\mathopen{}\mathclose{{\left[D(0,1)}}\right]^{2}$},$
@@
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left({\widehat{G}}_{m}^{(n)}-G_{{\tiny\mbox{mixt}}}}}\right)\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}$ $\sqrt{\pi_{0}\sigma^{2}}\mathbb{U}+\sqrt{\mathopen{}\mathclose{{\left(1-\pi_{0}}}\right)\sigma^{2}}\mathbb{V}_{G}+G_{{\tiny\mbox{mixt}}}^{\prime}\sqrt{1-\sigma^{2}}\mathbb{W}$
$=\pi_{0}\mathbb{Z}_{0}+\mathopen{}\mathclose{{\left(1-\pi_{0}}}\right)\mathbb{Z}_{1}\eqqcolon\mathbb{Z}\text{ on $D(0,1)$}.$
@@
Using Proposition 5.3 with the continuous functionals (45) gives us convergence in $D(0,1)$ of the asymptotic processes $(\mathrm{FDP}(t))_{t\in(0,1)}$ and $(\mathrm{TDP}(t))_{t\in(0,1)}$ and is left to the readers. This can be compared with the results in Genovese and Wasserman (2004) where they use independent and uniform $p$-values under the null. Here we have to deals with the dependence giving us the $\mathbb{W}$ term and the double asymptotic giving us the $\sigma^{2}$ term in Proposition 5.3. The same applies for the weighted case. However we are intesting to understand the asymptotic when the threshold is data-driven, and specifically the BH threshold.
By using the functional delta method theorem with the BH functionals (45) and (46) described above (see Neuvial, 2008, supplementary files from Delattre and Roquain, 2016 and Lemma S.2.2 from Kluger and Owen, 2024), we obtain the following result.
@@
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}({\widehat{G}}_{m}^{(n)})-\mathcal{T}_{\alpha}}}\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\frac{1}{\frac{1}{\alpha}-G_{{\tiny\mbox{mixt}}}^{\prime}(\mathcal{T}_{\alpha})}\mathbb{Z}({\mathcal{T}_{\alpha}});$
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left(\mathrm{FDP}_{m}^{(n)}\mathopen{}\mathclose{{\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}({\widehat{G}}_{m}^{(n)})}}\right)-\pi_{0}\alpha\frac{}{}}}\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\frac{\pi_{0}}{G_{{\tiny\mbox{mixt}}}(\mathcal{T}_{\alpha})}\mathbb{Z}_{0}({\mathcal{T}_{\alpha}});$
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left(\mathrm{TDP}_{m}^{(n)}\mathopen{}\mathclose{{\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}({\widehat{G}}_{m}^{(n)})}}\right)-G(\mathcal{T}_{\alpha})}}\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\frac{G^{\prime}(\mathcal{T}_{\alpha})}{\frac{1}{\alpha}-G_{{\tiny\mbox{mixt}}}^{\prime}(\mathcal{T}_{\alpha})}\mathbb{Z}({\mathcal{T}_{\alpha}})+\mathbb{Z}_{1}(\mathcal{T}_{\alpha}),$
@@
Proposition 5.4 is proved in Section B.3. As a sanity check, we see that Proposition 5.4 reduces to the result of Neuvial (2008) when $n/m$ tends to infinity ($\sigma^{2}=1$).
@@
$\begin{pmatrix}\mathbb{Z}_{0}\\ \mathbb{Z}_{1}\\ \mathbb{Z}\end{pmatrix}=\begin{pmatrix}\frac{\sigma}{\sqrt{\pi_{0}}}\mathbb{U}+\sqrt{1-\sigma^{2}}\mathbb{W}\\ \frac{\sigma}{\sqrt{1-\pi_{0}}}\mathbb{V}_{G}+G^{\prime}\sqrt{1-\sigma^{2}}\mathbb{W}\\ \sqrt{\pi_{0}\sigma^{2}}\mathbb{U}+\sqrt{\mathopen{}\mathclose{{\left(1-\pi_{0}}}\right)\sigma^{2}}\mathbb{V}_{G}+G_{{\tiny\mbox{mixt}}}^{\prime}\sqrt{1-\sigma^{2}}\mathbb{W}\end{pmatrix}$,
@@
Proving Theorem 4.2 is analogue to the proof of Theorem 4.1 above, but starting from weighted processes. For short, the full description of the proof is postponed to Section C.3. Therein, Propositions C.3 and C.4 are the analogues of Propositions 5.3 and 5.4, respectively.
@@ 6 Conclusion
In this paper we obtain the exact asymptotic distribution of $\mathrm{FCP}$ and $\mathrm{FDP}$ for conformal inference methods when both the sizes of the calibration sample and test sample grow simultaneously. Our theory covers both the prediction and novelty detection settings, including a potential distribution shift. Our results quantify exactly how the covariance process is affected by the dependence inherent to the conformal settings, that use the same calibration sample for all the test examples. First, we prove that the convergence rate $\sqrt{\tau_{n,m}}$ can be largely deteriorated when $n/m$ vanishes to zero (that is, $\sigma=0$). Otherwise, when $n$ and $m$ are of the same order, the convergence rate is the usual one ($m^{1/2}$), but the asymptotic covariance is affected. Nevertheless, when $n/m$ tends to infinity, the convergence is strictly the same as in the usual independent case. Interestingly, our results can be used to calibrate easily and accurately quantiles for controlling an error amount when performing conformal inference with large $n$ and $m$. We also quantify the effects of doing conformal inference while there is a distribution shift between the calibration sample and the test sample. We exhibit how this distribution shift acts on the asymptotic behaviour of the $\mathrm{FCP}$, by changing the mean and the variance, and how the correction with weighted conformal $p$-values impacts the asymptotic variance.
While our work paves the way for studying asymptotic convergences in conformal inferences, it leaves some open directions. For instance, in case of a distribution shift, the oracle weight function is mostly unknown, and is often estimated (Jin and Candès, 2023). Finding the exact asymptotic distribution for the processes using estimated weights is a very interesting and challenging problem for future investigations. In a finite sample setting, using the asymptotic expectation and variance as a proxy to design weight and score functions can provide very compelling new ideas and can unlock new training procedures.
{acks}
[Acknowledgments] I am grateful to Étienne Roquain for his precious advices and reviews throughout this work. I wish to thank Gilles Blanchard and Sylvain Delattre for their comments, and Magalie Fromont for fruitful discussions. I acknowledge the Emergence project MARS of Sorbonne Université, and grant ANR-23-CE40-0018-01 (BACKUP) of the French National Research Agency ANR.
@@ Appendix A Standardisation lemma
with $S^{\prime}_{k}=F_{\tiny\mbox{cal}}(S_{k})$ which are i.i.d. uniform and $T^{\prime}_{i}=F_{\tiny\mbox{cal}}(T_{i})$ which are i.i.d. $\sim{F_{\tiny\mbox{test,cal}}}$ (or either $\sim{F_{\tiny\mbox{$0$,cal}}}$ under the null or $\sim{F_{\tiny\mbox{test,cal}}}$ under the alternative, in the novelty detection setting). ∎
@@ Appendix B Proofs of auxiliary results
In this section, we prove Proposition 5.1, Propositions 5.3 and 5.4. Proofs for the weighted case are given in Section C.
@@
$\sqrt{m}\mathopen{}\mathclose{{\left({\widehat{F}}_{m,\tiny\mbox{test}}-F_{\tiny\mbox{test}}}}\right)\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\mathbb{U}\circ F_{\tiny\mbox{test}}\text{ on $D(\mathbb{R})$,}$
@@
$\sqrt{n}\mathopen{}\mathclose{{\left({\widehat{F}}_{\tiny\mbox{cal}}^{(n)}-F_{\tiny\mbox{cal}}}}\right)\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\mathbb{V}\circ F_{\tiny\mbox{cal}}\text{ on $D(\mathbb{R})$,}$
@@
$\sqrt{n}\mathopen{}\mathclose{{\left({({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}-F_{\tiny\mbox{cal}}^{-1}}}\right)\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}({F_{\tiny\mbox{cal}}^{-1}})^{\prime}\mathbb{V}\text{ on $D(0,1)$,}$ (49)
@@
$\sqrt{n}\mathopen{}\mathclose{{\left(F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}-F_{\tiny\mbox{test}}\circ F_{\tiny\mbox{cal}}^{-1}}}\right)\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}F_{\tiny\mbox{test}}^{\prime}\circ F_{\tiny\mbox{cal}}^{-1}({F_{\tiny\mbox{cal}}^{-1}})^{\prime}\mathbb{V}=\mathopen{}\mathclose{{\left(F_{\tiny\mbox{test}}\circ F_{\tiny\mbox{cal}}^{-1}}}\right)^{\prime}\mathbb{V}\text{ on $D(0,1)$.}$
Since for all $t\in(0,1)$, $G^{\prime}(t)=\mathopen{}\mathclose{{\left(F_{\tiny\mbox{test}}\circ F_{\tiny\mbox{cal}}^{-1}}}\right)^{\prime}(1-t)$ and since $\mathbb{V}$ is a standard Brownian bridge if and only if $(\mathbb{V}_{t})_{t}=(\mathbb{V}_{1-t})_{t}$ is a standard Brownian bridge we obtain the second term in (37) .
@@
$\begin{pmatrix}\sqrt{m}\mathopen{}\mathclose{{\left[{\widehat{F}}_{m,\tiny\mbox{test}}-F_{\tiny\mbox{test}}}}\right]\\ \sqrt{n}\mathopen{}\mathclose{{\left[F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ F_{\tiny\mbox{cal}}^{-1}(1-I)}}\right]\end{pmatrix}\overset{\mathcal{L}}{\rightarrow}\begin{pmatrix}\mathbb{U}\circ F_{\tiny\mbox{test}}\\ G^{\prime}\mathbb{V}\end{pmatrix}\text{ on $D(\mathbb{R})\times D(0,1)$},$
@@
$\begin{pmatrix}\sqrt{m}\mathopen{}\mathclose{{\left[{\widehat{F}}_{m,\tiny\mbox{test}}-F_{\tiny\mbox{test}}}}\right]\\ \sqrt{n}\mathopen{}\mathclose{{\left[F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ F_{\tiny\mbox{cal}}^{-1}(1-I)}}\right]\\ {({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)\end{pmatrix}\overset{\mathcal{L}}{\rightarrow}\begin{pmatrix}\mathbb{U}\circ F_{\tiny\mbox{test}}\\ G^{\prime}\mathbb{V}\\ F_{\tiny\mbox{cal}}^{-1}(1-I)\end{pmatrix}$
$\text{ on $D(\mathbb{R})\times D(0,1)\times\ell^{\infty}(0,1)$}.$
@@
$\begin{pmatrix}\sqrt{m}\mathopen{}\mathclose{{\left[{\widehat{F}}_{m,\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)}}\right]\\ \sqrt{n}\mathopen{}\mathclose{{\left[F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{(n)}})^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ F_{\tiny\mbox{cal}}^{-1}(1-I)}}\right]\end{pmatrix}\overset{\mathcal{L}}{\rightarrow}\begin{pmatrix}\mathbb{U}\mathopen{}\mathclose{{\left(F_{\tiny\mbox{test}}\circ F_{\tiny\mbox{cal}}^{-1}(1-I)}}\right)\\ G^{\prime}\mathbb{V}\end{pmatrix}$
$\text{ on $\mathopen{}\mathclose{{\left[D(0,1)}}\right]^{2}$}.$
Since $\mathbb{U}\mathopen{}\mathclose{{\left(F_{\tiny\mbox{test}}\circ F_{\tiny\mbox{cal}}^{-1}(1-I)}}\right)$ has the same distribution as $\mathbb{U}(G)$, we obtain (37).